### v1

@@ 0.3 Cavitation at a smooth defect-free surface
Representative snapshots of cavitation at surfaces with different contact angles are presented in Fig. 2 B. Fig. 2 C shows $1/{p^{*}_{\rm cav}}^{2}$ versus the negative pressure-ramp rate $-\dot{p}$. Each data point represents an average over 10 independent simulations. The solid lines are fits of the kinetic theory for $p_{\mathrm{2Dcav}}^{*}$ assuming pure surface cavitation given in Eq. 16, from which the values of $\kappa_{\textrm{2D}}(\theta)$ for each $\theta$ are obtained. For the lowest contact angle of 45, bubble cavitation occurs in the bulk. Therefore, this surface is not included in the analysis in Fig. 2 C. The resulting values for $\kappa_{\textrm{2D}}$ (symbols) are shown as a function of the surface contact angle $\theta$ in Fig. 2D. We also denote the characteristic length $\xi=\kappa_{\textrm{2D}}/\kappa_{\textrm{3D}}$ by the scale on the right, which represents the water slab height for which the surface cavitation attempt frequency $\kappa_{\textrm{2D}}$ equals the bulk attempt frequency $\kappa_{\textrm{3D}}$. It can be seen that $\xi$ exhibits a maximum around $\theta\approx 75^{\circ}$ and goes dramatically down with increasing contact angle from approximately $10$ to $10^{-2}$ nm. This suggests that a hydrophilic surface produces more cavitation nuclei in a given time than a hydrophobic surface, which is interesting since it contrasts with the pronounced density fluctuations at hydrophobic surfaces (45) (alternatively, it could reflect inaccuracies of the bulk and surface barrier free energy expressions). Clearly, this variation of $\kappa_{\textrm{2D}}$ is subdominant compared to the variation of the exponential barrier energy in Eq. 9, but nevertheless has to be included for a quantitative comparison of surface and bulk cavitation kinetics. For our subsequent analyses, we fit our simulation data for $\kappa_{\mathrm{2D}}$ in Fig. 2D to a function shown as a solid line, as detailed in the SI Appendix, section S1.
@@
Figure 3: (A) Cavitation pressure $p_{\mathrm{cav}}$ of a water-filled cubic container as a function of the contact angle $\theta$ of the defect-free inner walls, for a fixed waiting time of $\tau=1\,\mathrm{s}$, computed by numerically inverting Eq. 1 for $N_{\textrm{def}}=0$ (solid lines). Different colors represent different cube sizes $L$. Dotted lines show the 2D cavitation pressure according to Eq. 10, while dashed lines denote the 3D cavitation pressure based on Eq. 5. (B)  Crossover contact angle $\theta^{*}$, where $k_{\mathrm{2D}}=k_{\mathrm{3D}}$, as a function of container size $L$ for different waiting times $\tau$, according to Eq. 11.
@@
0.4 Kinetic competition between cavitation in bulk and at a defect-free surface
@@
In the absence of defects we set $N_{\mathrm{def}}=0$ in Eq. 1. The cavitation pressure $p_{\mathrm{cav}}$ cannot be derived analytically, therefore, we solve Eq. 1 numerically and plot $p_{\mathrm{cav}}$ in Fig. 3 A (solid lines) against the surface contact angle $\theta$ for different cube sizes $L$ and a fixed waiting time of $\tau=k_{\mathrm{tot}}^{-1}=1\,\mathrm{s}$. For hydrophilic surfaces with small $\theta$, $p_{\mathrm{cav}}$ is dominated by bulk cavitation (thus, $k_{\textrm{tot}}\approx k_{\textrm{3D}}$) and becomes independent of $\theta$, as described by Eq. 5 and represented by the horizontal dashed lines. For larger contact angles $\theta$ the cavitation pressure is well described by the surface prediction Eq. 10, shown as dotted lines, which indicates that cavitation for large $\theta$ predominantly occurs at surfaces.
@@
It is seen that cavitation shifts from the bulk (3D) to the surface (2D) pathway rather abruptly at a well-defined crossover contact angle $\theta^{*}$, which defines the onset of surface cavitation. Thus, $\theta^{*}$ is identified as the intersection of the bulk prediction (dashed lines) and the surface prediction (dotted lines) in Fig. 3 A. In SI Appendix, section S2, we present results for a range of different waiting times $\tau$ from $10^{-3}$ to $10^{3}$ s, showing the same sharp transition between the surface-dominated and the bulk-dominated regimes. By equating Eq. 5 and Eq. 10 we obtain an exact expression for the critical system size $L^{*}$ at which the transition takes place
@@
see SI Appendix, section S3 for the derivation. The critical contact angle depends as a quartic square root on the logarithm of $L$ and $\tau$, which is an extremely weakly varying function.
We conclude that surface cavitation occurs if the contact angle of the confining surfaces exceeds the threshold value $\theta^{*}$, which is largely independent of system size and waiting time. Although surfaces with contact angles below the threshold, i.e. for $\theta<\theta^{*}$, do have a lower cavitation free energy barrier compared to bulk cavitation, as predicted by Eq. 7, they exhibit a smaller cavitation rate due to the different attempt frequencies at a surface and in bulk, as quantified by the length scale $\xi$. This renders surface cavitation on hydrophilic surfaces less likely than bulk cavitation.
0.5 Cavitation at surface defects
Real-world surfaces are rich in surface defects such as vacancies, disordered regions, variations in material composition and adsorbed organic or inorganic species (46). We model defects as circular patches with radius $R_{\mathrm{def}}$ and contact angle $\theta_{\mathrm{def}}$, as illustrated in Fig. 4 A. We assume the defects to be larger than the critical bubble base radius $a^{*}$ and much smaller than the system size $L$, $a^{*}\ll R_{\mathrm{def}}\ll L$. With these assumptions, the single defect cavitation rate in Eq. 1 is given in analogy to Eq. 9 as
@@ 0.5 Cavitation at surface defects
with the defect surface area given by $A_{\mathrm{def}}=\pi R_{\mathrm{def}}^{2}$ and the attempt frequency density $\kappa_{\mathrm{2D}}$ and the free energy barrier $G^{*}_{\mathrm{2D}}$ determined by the defect contact angle $\theta_{\mathrm{def}}$. The defect cavitation pressure follows from Eq. 13 by using Eq. 7 and $\tau=1/(N_{\mathrm{def}}k_{\mathrm{def}})$ as
@@
The total cavitation rate in Eq. 1 describes the kinetic competition between cavitation in bulk, at surfaces and at defects. The cavitation pressure resulting from this competition is determined by solving Eq. 1 in conjunction with Eqs. 4, 9, 13 for a given set of parameters $L$, $\tau$, $\theta$, $N_{\mathrm{def}}$ and $\theta_{\mathrm{def}}$.
@@
Figure 4: (A) Illustration of a cavitation bubble on a circular surface defect with radius $R_{\textrm{def}}$ and contact angle $\theta_{\textrm{def}}$ on a flat surface of contact angle $\theta$. (B) Cavitation pressure of a water-filled container with a single surface defect $N_{\textrm{def}}=1$ as a function of $R_{\textrm{def}}$ for different $\theta_{\textrm{def}}$. Solid lines are obtained by numerically solving Eq. 1 for a waiting time of $\tau=1\,\mathrm{s}$ and a container size of $L=1\,\mathrm{cm}$ (volume of $1\,\mathrm{mL}$). Dotted lines show the limiting pressure for a large defect, given by Eq. 14, whereas the dashed line shows the bulk cavitation pressure Eq. 5. (C) Crossover defect radius as a function of container size $L$ for a fixed waiting time of $\tau=1\,\mathrm{s}$ from Eq. 15. (D) Simulation snapshots for box expansions of $L_{z}/L_{z0}=3/2$ after cavitation in systems with a surface contact angle of $97^{\circ}$ and cylindrical hydrophobic pits, from pressure-ramp simulations at a rate of $\dot{p}=-5\,\mathrm{MPa/ns}$. The pits have radii of $R_{\textrm{def}}\approx 1$ nm and $2\leavevmode\nobreak\ \mathrm{nm}$.
@@
For the following discussion, we assume that the surface is sufficiently hydrophilic so that $k_{\mathrm{2D}}\ll k_{\mathrm{3D}}$ and cavitation at the defect-free surface parts can be neglected. The solid lines in Fig. 4B show the cavitation pressure in a 1 mL cube of water ($L=1$ cm) in the presence of a single surface defect, $N_{\mathrm{def}}=1$, as a function of the defect radius $R_{\mathrm{def}}$ for different defect contact angles $\theta_{\textrm{def}}$. For a given $\theta_{\textrm{def}}$ a well-defined crossover radius $R^{*}_{\mathrm{def}}$ emerges, at which the system transitions from bulk cavitation for small $R_{\mathrm{def}}$, where the bulk prediction Eq. 5 (dashed line) is valid, to defect cavitation for large $R_{\mathrm{def}}$, where the prediction in Eq. 14 (dotted lines) is valid. The crossover radius $R^{*}_{\mathrm{def}}$ grows as $\theta_{\textrm{def}}$ approaches $\theta^{*}$ from above. Conversely, for increasing $\theta_{\textrm{def}}$ the crossover radius $R^{*}_{\mathrm{def}}$ approaches sub-nanometer values (where our continuum model will eventually break down). The results in Fig. 4B show that cavitation is dominated by defects with a radius larger than a critical value $R^{*}_{\mathrm{def}}$ and occurs at a rate $N_{\textrm{def}}k_{\textrm{def}}$; even a single hydrophobic defect can dominate the cavitation kinetics in macroscopic systems.
@@
The crossover radius $R^{*}_{\mathrm{def}}$ between the cavitation pressures for the bulk and defect pathways follows from equating Eqs. 5 and 14 as
@@
In Fig. 4 C we show $R_{\mathrm{def}}^{*}$ as a function of system size $L$ for a waiting time $\tau=1\,\mathrm{s}$ and a single defect $N_{\mathrm{def}}=1$ for three different defect contact angles. As the system size increases, a larger defect is required to induce defect cavitation. Conversely, for larger defect contact angle $\theta_{\textrm{def}}$ the crossover radius $R_{\textrm{def}}^{*}$ decreases and reaches a nanometer even for macroscopically large systems. For a surface with many defects, $N_{\mathrm{def}}>1$, the required defect radius $R_{\mathrm{def}}^{*}$ is from Eq. 15 predicted to go further down.
@@
Finally, we check our theoretical predictions by explicit simulations of a system featuring one defect. We choose a mildly hydrophobic surface with a contact angle of $\theta=97^{\circ}$ on which we create one circular defect by removing surface molecules within a radius of $R_{\mathrm{def}}$ on one monolayer, thereby forming a hydrophobic pit with a depth of 1 nm. Water does not enter the pit because of its hydrophobic interior, creating a vapor bubble, corresponding to a defect contact angle of $\theta_{\textrm{def}}\approx 180^{\circ}$. To avoid cavitation at the opposite monolayer, we make it hydrophilic with a contact angle of $42^{\circ}$. Using the pressure ramp protocol, we consistently observe bubble formation at the hydrophobic pit, as shown in Fig. 4 D, despite the rest of the surface being mildly hydrophobic. Using a pressure rate of $\dot{p}=-5\,\mathrm{MPa/ns}$, a defect-free surface induces surface cavitation at $p^{*}_{\mathrm{cav}}=-95\pm 1$ MPa, as seen in Fig. 2 C. In contrast, the systems with surface defects of 1 and 2 nm radii cavitate already at $p_{\mathrm{cav}}^{*}=-66$ and $-45\pm 2$ MPa, respectively, in qualitative agreement with our predictions in Fig. 4 B.
Conclusion

### v2

@@ III.3 Cavitation at a smooth defect-free surface
Representative snapshots of cavitation for $L_{z}/L_{z}^{0}=3/2$ at surfaces with different contact angles are presented in Fig. 3 B. We also provide trajectory visualizations for $\theta=45^{\circ}$ (upper left panel) (multimedia available online) and $\theta=135^{\circ}$ (lower right panel) (multimedia available online). Fig. 3 C shows $1/{p^{*}_{\rm cav}}^{2}$ versus the negative pressure-ramp rate $-\dot{p}$. Each data point represents an average over 10 independent simulations. The solid lines are fits of the kinetic theory (Eq. 1 and using $G^{*}=G^{*}_{\mathrm{2D}}$), assuming pure surface cavitation which yield the value of $\kappa_{\textrm{2D}}(\theta)$ for each $\theta$. For the lowest contact angle of 45, bubble cavitation occurs in the bulk. Therefore, this surface is not included in the analysis in Fig. 3 C. The resulting values for $\kappa_{\textrm{2D}}$ (symbols) are shown as a function of the surface contact angle $\theta$ in Fig. 3D. We also denote the characteristic length $\xi=\kappa_{\textrm{2D}}/\kappa_{\textrm{3D}}$ by the scale on the right, which represents the water slab height for which the surface cavitation attempt frequency $\kappa_{\textrm{2D}}$ equals the bulk attempt frequency $\kappa_{\textrm{3D}}$. It can be seen that $\xi$ exhibits a maximum around $\theta\approx 75^{\circ}$ and goes dramatically down with increasing contact angle from approximately $10$ to $10^{-2}$ nm. This suggests that a hydrophilic surface produces more cavitation nuclei in a given time than a hydrophobic surface, which is interesting since it contrasts with the pronounced density fluctuations at hydrophobic surfaces Rego and Patel (2022) (alternatively, it could reflect inaccuracies of the bulk and surface-barrier-free energy expressions). Clearly, this variation of $\kappa_{\textrm{2D}}$ is subdominant compared to the variation of the exponential barrier energy in Eq. 11, but nevertheless has to be included for a quantitative comparison of surface and bulk cavitation kinetics. For our subsequent analyses, we fit our simulation data for the logarithm of $\kappa_{\mathrm{2D}}$ in Fig. 3D to a polynomial shown as a solid line, as detailed in the Supplementary Material, section S3. While polynomial fits are suitable for interpolation, they are less reliable for extrapolation. We therefore trust the polynomial fit only down to a constant angle extrapolation intervall $60^{\circ}-\theta^{*}=8^{\circ}$ which is half the angle distance between the two lower data points $76^{\circ}-60^{\circ}=16^{\circ}$. Therefore the fit for $\theta<\theta^{*}=52^{\circ}$ is shown as a broken line.
@@
Figure 4: (A) Cavitation pressure $p_{\mathrm{cav}}$ of a water-filled cubic container as a function of the contact angle $\theta$ of the defect-free inner walls, for a fixed waiting time of $\tau=1\,\mathrm{s}$, computed by numerically inverting Eq. 3 for $N_{\textrm{def}}=0$ (solid lines). Different colors represent different cube sizes $L$. Dotted lines show the 2D cavitation pressure according to Eq. 12, while dashed horizontal lines denote the 3D cavitation pressure based on Eq. 7. (B)  Crossover contact angle $\theta^{*}$, where $k_{\mathrm{2D}}=k_{\mathrm{3D}}$, as a function of container size $L$ for different waiting times $\tau$, according to Eq. 13.
@@
III.4 Kinetic competition between cavitation in bulk and at a defect-free surface
@@
In the absence of defects, we set $N_{\mathrm{def}}=0$ in Eq. 3. The cavitation pressure $p_{\mathrm{cav}}$ cannot be derived analytically, therefore, we solve Eq. 3 numerically and plot $p_{\mathrm{cav}}$ in Fig. 4 A (solid lines) against the surface contact angle $\theta$ for different cube sizes $L$ and a fixed waiting time of $\tau=k_{\mathrm{tot}}^{-1}=1\,\mathrm{s}$. For hydrophilic surfaces with small $\theta$, $p_{\mathrm{cav}}$ is dominated by bulk cavitation (thus, $k_{\textrm{tot}}\approx k_{\textrm{3D}}$) and becomes independent of $\theta$, as described by Eq. 7 and represented by the horizontal dashed lines. For larger contact angles $\theta$, the cavitation pressure is well described by the surface prediction (Eq. 12), shown as dotted lines, which indicates that cavitation for large $\theta$ predominantly occurs at surfaces.
@@
It is seen that cavitation shifts from the bulk (3D) to the surface (2D) pathway rather abruptly at a distinct crossover contact angle $\theta^{*}$, which marks the onset of surface cavitation. Thus, $\theta^{*}$ is identified as the intersection of the bulk prediction (dashed lines) and the surface prediction (dotted lines) in Fig. 4 A. In Supplementary Material, section S4, we present results for a range of different waiting times $\tau$ from $10^{-3}$ to $10^{3}$ s, showing the same sharp transition between the surface-dominated and the bulk-dominated regimes. By equating Eq. 7 and Eq. 12, we obtain an exact expression for the critical system size $L^{*}$ at which the transition takes place
@@
see Supplementary Material, section S5 for the derivation. The critical contact angle depends as a fourth root on the logarithm of $L$ and $\tau$, which is an extremely weakly varying function.
We conclude that surface cavitation occurs if the contact angle of the confining surfaces exceeds the threshold value $\theta^{*}$, which is largely independent of system size and waiting time. Although surfaces with contact angles below the threshold (i.e., for $\theta<\theta^{*}$) do have a lower cavitation free energy barrier compared to bulk cavitation, as predicted by Eq. 9, they exhibit a smaller cavitation rate due to the different attempt frequencies at a surface and in bulk, as quantified by the length scale $\xi$. This renders surface cavitation on hydrophilic surfaces less likely than bulk cavitation.
III.5 Cavitation at surface defects
Real-world surfaces are rich in surface defects such as vacancies, disordered regions, variations in material composition and adsorbed organic or inorganic species Siretanu, van den Ende, and Mugele (2016). We model defects as circular patches with radius $R_{\mathrm{def}}$ and contact angle $\theta_{\mathrm{def}}$, as illustrated in Fig. 5 A. We assume the defects to be larger than the critical bubble base radius $a^{*}$ and much smaller than the system size $L$, $a^{*}\ll R_{\mathrm{def}}\ll L$. With these assumptions, the single defect cavitation rate in Eq. 3 is given in analogy to Eq. 11 as
@@ III.5 Cavitation at surface defects
with the defect surface area given by $A_{\mathrm{def}}=\pi R_{\mathrm{def}}^{2}$ and the attempt frequency density $\kappa_{\mathrm{2D}}$ and the free energy barrier $G^{*}_{\mathrm{2D}}$ determined by the defect contact angle $\theta_{\mathrm{def}}$. The defect cavitation pressure follows from $\tau^{-1}=N_{\mathrm{def}}k_{\mathrm{def}}$ by using Eq. 15 and Eq. 9 as
@@
The total cavitation rate in Eq. 3 describes the kinetic competition between cavitation in bulk, at surfaces and at defects. The cavitation pressure resulting from this competition is determined by solving Eq. 3 in conjunction with Eqs. 6, 11, 15 for a given set of parameters $L$, $\tau$, $\theta$, $N_{\mathrm{def}}$, $\theta_{\mathrm{def}}$ and $R_{\mathrm{def}}$.
@@
Figure 5: (A) Illustration of a cavitation bubble on a circular surface defect with radius $R_{\textrm{def}}$ and contact angle $\theta_{\textrm{def}}$ on a flat surface of contact angle $\theta$. (B) Cavitation pressure of a water-filled container with a single surface defect $N_{\textrm{def}}=1$ as a function of $R_{\textrm{def}}$ for different $\theta_{\textrm{def}}$. Solid lines are obtained by numerically solving Eq. 3 for a waiting time of $\tau=1\,\mathrm{s}$ and a container size of $L=1\,\mathrm{cm}$ (volume of $1\,\mathrm{mL}$). Dotted lines show the limiting pressure for a large defect, given by Eq. 16, whereas the horizontal dashed line shows the bulk cavitation pressure Eq. 7. (C) Crossover defect radius $R_{\mathrm{def}}^{*}$ as a function of container size $L$ for a fixed waiting time of $\tau=1\,\mathrm{s}$ from Eq. 17. (D) Simulation snapshots for box expansions of $L_{z}/L_{z}^{0}=3/2$ after cavitation in systems with a surface contact angle of $97^{\circ}$ and cylindrical hydrophobic pits, from pressure-ramp simulations at a rate of $\dot{p}=-5\,\mathrm{MPa/ns}$. The pits have radii of $R_{\textrm{def}}\approx 1$ nm (upper panel) (multimedia available online) and $2\leavevmode\nobreak\ \mathrm{nm}$ (lower panel) (multimedia available online).
@@
For the following discussion, we assume that the surface is sufficiently hydrophilic so that $k_{\mathrm{2D}}\ll k_{\mathrm{3D}}$ and cavitation at the defect-free surface parts can be neglected. The solid lines in Fig. 5B show the cavitation pressure in a 1 mL cube of water ($L=1$ cm) in the presence of a single surface defect, $N_{\mathrm{def}}=1$, as a function of the defect radius $R_{\mathrm{def}}$ for different defect contact angles $\theta_{\textrm{def}}$. For a given $\theta_{\textrm{def}}$, a well-defined crossover radius $R^{*}_{\mathrm{def}}$ emerges at which the system transitions from bulk cavitation for small $R_{\mathrm{def}}$, where the bulk prediction Eq. 7 (dashed line) is valid, to defect cavitation for large $R_{\mathrm{def}}$, where the prediction in Eq. 16 (dotted lines) is valid. The crossover radius $R^{*}_{\mathrm{def}}$ grows as $\theta_{\textrm{def}}$ approaches $\theta^{*}$ from above. Conversely, for increasing $\theta_{\textrm{def}}$ the crossover radius $R^{*}_{\mathrm{def}}$ approaches sub-nanometer values (where our continuum model will eventually break down). The results in Fig. 5B show that cavitation is dominated by defects with a radius larger than a critical value $R^{*}_{\mathrm{def}}$ and occurs at a rate $N_{\textrm{def}}k_{\textrm{def}}$; even a single hydrophobic defect can dominate the cavitation kinetics in macroscopic systems.
@@
The crossover radius $R^{*}_{\mathrm{def}}$ between the cavitation pressures for the bulk and defect pathways follows from equating Eqs. 7 and 16 as
@@
In Fig. 5 C we show $R_{\mathrm{def}}^{*}$ as a function of system size $L$ for a waiting time of $\tau=1\,\mathrm{s}$ and a single defect ($N_{\mathrm{def}}=1$) for four different defect contact angles. As the system size increases, a larger defect is required to induce defect cavitation. Conversely, for larger defect contact angle $\theta_{\textrm{def}}$, the crossover radius $R_{\textrm{def}}^{*}$ decreases and reaches the nanometer scale even for macroscopically large systems. For a surface with multiple defects, $N_{\mathrm{def}}>1$, the required defect radius $R_{\mathrm{def}}^{*}$ is expected to decrease proportionally to $\sim N_{\mathrm{def}}^{-1/2}$, as predicted by Eq. 17.
@@
Finally, we check our theoretical predictions by explicit simulations of a system featuring one defect. We choose a mildly hydrophobic surface with a contact angle of $\theta=97^{\circ}$ on which we create one circular defect by removing surface molecules within a radius of $R_{\mathrm{def}}$ on one monolayer, thereby forming a hydrophobic pit with a depth of 1 nm. Water does not enter the pit because of its hydrophobic interior, creating a vapor bubble, corresponding to a defect contact angle of $\theta_{\textrm{def}}\approx 180^{\circ}$. To avoid cavitation at the opposite monolayer, we make it hydrophilic with a contact angle of $42^{\circ}$. Using the pressure ramp protocol, we consistently observe bubble formation at the hydrophobic pit, as shown in the snapshots in Fig. 5 D, despite the rest of the surface being mildly hydrophobic. Trajectory visualizations for $R_{\mathrm{def}}=1$ nm (upper panel) (multimedia available online) and $R_{\mathrm{def}}=2$ nm (lower panel) (multimedia available online) defects are provided as well. Using a pressure rate of $\dot{p}=-5\,\mathrm{MPa/ns}$, a defect-free surface induces surface cavitation at $p^{*}_{\mathrm{cav}}=-95\pm 1$ MPa, as seen in Fig. 3 C. In contrast, the systems with surface defects of 1 and 2 nm radii cavitate already at $p_{\mathrm{cav}}^{*}=-66$ and $-45\pm 2$ MPa, respectively. These results qualitatively agree with our predictions in Fig. 5 B that even nanometer-sized surface defects can act as cavitation nuclei. The influence of defects on cavitation is studied in more detail in our follow-up study Šako et al. (2024).
IV Conclusion